Hanna:

If the combined area of the white and orange regions is 100π, what is the area of ONLY the orange region? A) 16π B) 36π C) 64π D) 96π

MARC:

Formula for area of a circle \(=\pi~r^{2}\)

MARC:

Total combined area \(=100~\pi\)

MARC:

\(\pi~6^{2}+\pi~r^{2}=100\pi\)

MARC:

\(36\pi+\pi~r^{2}=100\pi\)

MARC:

Factorise pi \(\pi(36+r^{2})=100\pi\)

MARC:

Divide both sides by pi \(36+r^{2}=100\)

MARC:

Subtract both sides by 36 \(r^{2}=64\)

MARC:

Square root both sides \(r=\pm8\)

MARC:

i realised that i give the answer... xD Sorry...

MARC:

okay,can u tell me why r=-8 cannot be accepted but r=8 can be accepted? @Hanna

Hanna:

because its a negative?

MARC:

yep,that's right! area can't hv negative value... :)

Hanna:

okay thx <3

MARC:

Np ^_^